OpenStudy (anonymous):

A pair of 388.8 N vectors are perpendicular. What is the magnitude of their resultant? Answer in units of N

OpenStudy (aaronq):

 \(\sf \Large c^2=a^2+b^2\) \(\sf \Large c=\sqrt{(388.8~N)^2+(388.8~N)^2}=549.846\approx 549.9~N\)

OpenStudy (anonymous):

thank you